A projectile is launched vertically from the surface of the Moon with an initial speed of 1460 m/s. At what altitude is the projectile's speed one-half its initial value?
Answer #1

Given,

The initial speed of the projectile, u\ =\ 1460\ m/s

The final speed of the projectile, v\ =\ \dfrac{u}{2}

  =\ \dfrac{1460}{2}\ m/s

  =\ 730\ m/s

Acceleration due to gravity on the surface of the moon, g\ =\ 1.625\ m/s

Since the projectile is going up so there will be de-acceleration in the projectile, hence from the equation of motion,

v^2\ =\ u^2-2gh

where h is the height reached by the projectile.4

h\ =\ \dfrac{u^2-v^2}{2g}

  =\ \dfrac{(1460)^2-(730)^2}{2\times 1.625}

  =\ 491907.69\ m

  =\ 491.907\ km

Hence, the altitude at which the projectile's speed is one-half of its initial value is 491.907 km.